Flounder provides environmentally friendly lawn services for homeowners. Its operating costs are as follows. Depreciation $1,400 per month Advertising $200 per month Insurance $7,640 per month Weed and feed materials $12 per lawn Direct labor $44 per lawn Fuel $2 per lawn Flounder charges $100 per treatment for the average single-family lawn. Incorrect answer. Your answer is incorrect. Determine the company’s break-even point in number of lawns serviced per month. Break-even point Entry field with incorrect answer lawns SHOW SOLUTION SHOW ANSWER LINK TO TEXT LINK TO TEXT Incorrect answer. Your answer is incorrect. Determine the company’s break-even point in dollars. (Round answer to 0 decimal places, e.g. 1,225.) Break-even point $Entry field with incorrect answer
Answer #1

Variable cost per lawn = 12+44+2 = 58

Fixed cost = 1400+200+7640 = 9240

a) Break even point = Fixed cost/Contribution margin per lawn = 9240/(100-58) = 220 Lawn

b) Break even sales = 220*100 = $22000